OpenStudy (blondisaurus):

Given that tan Ѳ = sin Ѳ/ cos Ѳ , show that 1+ cot^2 Ѳ = csc^2Ѳ is a true trigonometric identity

OpenStudy (anonymous):

\(\tan(\theta) = \frac{\sin(\theta)}{\cos(\theta)} \implies \cot(\theta) = \frac{\cos(\theta)}{\sin(\theta)}\), so \(\cot^2(\theta) = \frac{\cos^2(\theta)}{\sin^2(\theta)}\), Adding one: \(\cot^2(\theta) + 1 = \frac{\cos^2(\theta)}{\sin^2(\theta)} + \frac{\sin^2(\theta)}{\sin^2(\theta)}\) \(=\frac{\cos^2(\theta) + \sin^2(\theta)}{\sin^2(\theta)}\) \(=\frac{1}{\sin^2(\theta)}\) \(= \csc^2(\theta)\)

OpenStudy (blondisaurus):

I get it completely. Thank you very much.

OpenStudy (anonymous):

No problem :).